1. If you had 1.50 L of 2.00-M solution, how much water would you need to add to obtain a 1.750-M solution?


2. Describe you would make the following solutions for 2.75L Of 0.0120 M NaOH
1
Expert's answer
2020-04-20T15:06:39-0400

1. Cm=n/V

n=Cm x V = 1.5 x 2 = 3 moles.

V= n/Cm = 3/1.75 = 1.715 liter.

1.715 - 1.5 = 0.215 L.

2. 2.75 x 0.012 = 0.033 moles of NaOH and 2.75 L water